OpenStudy (anonymous):

(3m^3 -1/2 y)(3m^3 -1/2 y) is tehre anyone who can take me step by step through this usign these numbers plese i realy need to get this assignment finished but im not understanding this i have to use the foil method

OpenStudy (klimenkov):

I can help you. But I really don't understand what you want.

OpenStudy (klimenkov):

Ah. Ok. Are you here?

OpenStudy (anonymous):

yes im here i have to simplify this usint the foil method and i dont understand it and if someone could show it to me step bye step i would get it and i could solve the rest on my own but i need some one to do it step bye step just the way im sendign this each step on a line of its own i have dyslexia and can be hard to read the problem cose up

OpenStudy (zehanz):

YOu need to FOIL\[(3m^3-\frac{ 1 }{ 2 }y)(3m^3-\frac{ 1 }{ 2 }y)\]Plain-vanilla FOIL goes like this:\[(a+b)(c+d)=ac(F)+ad(O)+bc(I)+bd(L)\]so try to do that with your numbers:\[F: 3m^3 \cdot 3m^3=...\]\[O:3m^3 \cdot -\frac{ 1 }{ 2 }y=...\]\[I:-\frac{ 1 }{ 2 }y \cdot 3m^3=...\]\[L:-\frac{ 1 }{ 2 }y \cdot -\frac{ 1 }{ 2 }y=...\]Then try piecieng the bits together where possible!

OpenStudy (zehanz):

Maybe this also helps: F =First - multiply the first number in the left pair of ( ) with the first one in the right O = Outer I = Inner L = Last

hero (hero):

What I showed WAS FOIL. The student just simply does not get it and spends more time complaining rather than trying to understand.

OpenStudy (zehanz):

@Hero: I missed that completely, I'm afraid...

OpenStudy (anonymous):

actualy i did get it and i figured out the right answere on my own @hero so thank you very much and now i have finished my assignment bc all five probs are done and i got a 90 so thank you